Southern Rim Parts estimates its manufacturing overhead to be $409,500 and its direct labor costs to be $910,000 for year 1.Southern Rim Parts estimates its manufacturing overhead to be $409,500 and its direct labor costs to be $910,000 for year 1.
Answer #1

Predetermined Overhead rate = Estimated overhead/Direct labor cost

= 409,500/910,000 = 45%

Job 301 (50,000*45%) 22,500
Job 302 (75,000*45%) 33,750
Job 303 (100,000*45%) 45,000

Applied overhead = 849,000*45% = 382,050

Underapplied overhead = 91,950